SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

   
Prev Next


Date: March 25, 2008
Shift Taker: Chukwudi K. Clarke
-----------
General Features:
(theta,phi)

Warm crystals in No cuts hitmap (17,43), (42,84), (52,63), (57,63), (56,64)
Warm crystals in 0.01 GeV hitmap (17,43), (30,45), (32,46), (36,64)
Warm crystals in 0.1 GeV hitmap (30,45), (32,46)
Warm crystals in 1 GeV hitmap      

Cool crystal  in hitmap          
Cool crystal  in 0.1 GeV hitmap (29,84) 
Cool crystal  in 1 GeV hitmap

0080960 Ok
0080961 Ok
0080962 Ok
0080963 Ok
0080964 Ok
0080965 Cosmics
0080966 >10^-3 TC damage in OutofSynch OutOfOrder FLINKA_TimeOut
FLINKB_TimeOut FLINKC_TimeOut ROM 0-5 1-3 3-0 and Fibres A-C.
0080967 10^-3 TC damage in OutofSynch OutOfOrder ROM 0-5 and Fibres A-C.
0080968 Ok
0080969 Ok
0080970 Ok
0080971 Ok
0080972 Ok
0080973 Ok
0080974 Ok
0080975 Ok
0080976 Ok
0080977 Horizontal line in Digi Timing plot @log(E)=0.4. Warm crystal @
(3,57) in Ecut=0.01 to 1GeV hit maps.
0080978 Ok
0080979 Ok
0080980 Ok
0080981 Ok
0080982 >10^-3 TC damage in OutofSynch OutOfOrder ROM 1-3 5-0 and Fibres A-C.
0080983 Ok
0080984 Ok
0080985 Ok
0080986 Ok
0080987 >10^-3 TC damage in OutofSynch OutOfOrder ROM 0-5 and Fibres A-C.

Crystal Spectra:
NTR
Lightpulser:
4 channels dead, 0 channels out of time. Last updated on: 25-March 2008 06:30 PST


 

  Links

  • Shift Report Index
  • Back to the Shifter instructions